OpenStudy (anonymous):

Simplify by dividing - 5/8 by -3/4

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

just a second

OpenStudy (texaschic101):

when dividing fractions, flip what you are dividing by, then multiply (-5/8) / (-3/4) = -5/8 * -4/3 = (neg times a neg is a positive) 20/24 = 5/6

OpenStudy (anonymous):

-5/6 ?

OpenStudy (texaschic101):

make sure you slip what you are dividing by and not the other number

OpenStudy (texaschic101):

* flip

OpenStudy (anonymous):

the multiple choice is 20/24, - 20/24, - 15/32, and 15/32

OpenStudy (texaschic101):

20/24...they didn't reduce

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

no problem :)